OpenStudy (anonymous):

28 beats in 10 seconds is how many beats per minute?

OpenStudy (anonymous):

168 beats per minute

OpenStudy (anonymous):

(Show work).

OpenStudy (anonymous):

28 beats per 10 second * 60 seconds per minute = 168 beats per minute

jhonyy9 (jhonyy9):

6*28= ?

OpenStudy (anonymous):

28 --10 second x ----60 secondss cross multiply \[\LARGE 60\cdot 28=10x \Longrightarrow x=6\cdot 28\]

OpenStudy (anonymous):

I suck at this drawing thing, so I did it in paint, but the method is the same regardless of what the numbers are, and you can apply the same concept to other equations, like minutes in an hour, or miles per hour, or anything per anything. As long as you have the conversion amount for the multiplication.

OpenStudy (anonymous):

IT WOULD BE 60 TIMES 28 BECAUSE BECAUSE THERES 60 SECONDS IN A MINTUE

OpenStudy (anonymous):

Caps lock much? Especially on a question that's already been answered.